A force of 0.004 N acts upon a body and speed of the body changes from 0.30m/sec to 0.10m/sec in passing through a certain distance . find the distance if mass of body is 1 kg.

Grade:11

1 Answers

sonika p
6 years ago
F = 0.004N
acceleration,a = F/m = 0.004m/s2
u = 0.3m/s
v = 0.1m/s
s = (v2 – u2 )/2a = (0.12-0.32)/2*0.004 = 10m
distance travelled by the body  is 10m